# Valid F-screening in linear regression

### v1

@@ Abstract
Suppose that a data analyst wishes to report the results of a least squares linear regression only if the overall null hypothesis—$H_{0}^{1:p}:\beta_{1}=\beta_{2}=\ldots=\beta_{p}=0$—is rejected. This practice, which we refer to as F-screening (since the overall null hypothesis is typically tested using an $F$-statistic), is in fact common practice across a number of applied fields. Unfortunately, it poses a problem: standard guarantees for the inferential outputs of linear regression, such as Type 1 error control of hypothesis tests and nominal coverage of confidence intervals, hold unconditionally, but fail to hold conditional on rejection of the overall null hypothesis. In this paper, we develop an inferential toolbox for the coefficients in a least squares model that are valid conditional on rejection of the overall null hypothesis. We develop selective p-values that lead to tests that control the selective Type 1 error, i.e., the Type 1 error conditional on having rejected the overall null hypothesis. Furthermore, they can be computed without access to the raw data, i.e., using only the standard outputs of a least squares linear regression, and therefore are suitable for use in a retrospective analysis of a published study. We also develop confidence intervals that attain nominal selective coverage, and point estimates that account for having rejected the overall null hypothesis. We show empirically that our selective procedure is preferable to an alternative approach that relies on sample splitting, and we demonstrate its performance via re-analysis of two datasets from the biomedical literature.
@@ 1 Introduction
at level $\alpha_{0}$ in order to determine whether the model in (1) is “interesting,” i.e., whether it merits further study. If they fail to reject $H_{0}^{1:p}$, then (1) receives no further consideration (and the fact that (1) was considered may not even be reported). On the other hand, if (and only if) they reject $H_{0}^{1:p}$, they then proceed to test $H_{0}^{M}:\beta_{M}=0$ for fixed $M\subseteq\left\{1,\ldots,p\right\}$, where $\beta_{M}:=\{\beta_{j}:j\in M\}.$ Because $H_{0}^{1:p}$ is typically tested by means of an $F$-statistic, we refer to this procedure—summarized in Box 1—as F-screening. If hypothesis tests are conducted in Step 2 of Box 1, then these are sometimes referred to as post hoc tests.
@@
[!h] Box 1: F-screening in the model (1). Step 1: Test $H_{0}^{1:p}$ as given in (2) at level $\alpha_{0}$, using an $F$-test. Step 2: Only if $H_{0}^{1:p}$ is rejected, conduct inference on $\beta_{M}$ for fixed $M\subseteq\{1,\ldots,p\}$.
@@
To see that conducting “standard” inference in Step 2 of F-screening will not provide standard statistical guarantees, consider the (extreme) setting where $p=1$ and $\beta_{1}=0$. In this case, $H_{0}^{1:p}\equiv H_{0}^{1}$. Suppose we carry out F-screening as in Box 1, using the “standard” $t$-test for $H_{0}^{1}$ in Step 2. Then the tests in Steps 1 and 2 are identical (if conducted at the same level), and so we will report a result in Step 2 only if $H_{0}^{1}$ is rejected. Thus, under the null hypothesis, the Type 1 error rate (among reported results) equals one! The left plot in Fig. 4 illustrates the inflation of Type 1 error that results from conducting “standard” inference in Step 2 in a setting with $p=10$ and $n=100$.
@@
As pointed out by Olshen (1973), conducting a test of $H_{0}^{M}$ that accounts for the fact that $H_{0}^{1:p}$ was rejected is entirely distinct from the task of correcting multiple tests of the form $H_{0}^{M}:\beta_{M}=0$ (for multiple sets $M$) for multiple comparisons. The latter has been very well-explored in the literature: see Dunn (1961); Šidák (1967); Tukey (1949); Scheffé (1953). For instance, if our interest in Step 2 of Box 1 lies in $H_{0}^{j}:\beta_{j}=0$ for $j\in\{1,\ldots,p\}$, then we could consider testing each null hypothesis using a standard test, and applying a Bonferroni correction to account for the fact that $p$ tests were conducted. However, this will not address the fact that F-screening was performed: namely, that we test $H_{0}^{j}$, for a fixed index $j$, if and only if we have rejected $H_{0}^{1:p}$. We illustrate the inadequacy of this approach numerically in Appendix A.
@@
Our goal of conducting inference on $\beta_{M}$ that is valid conditional on the fact that we rejected $H_{0}^{1:p}$ falls under the umbrella of selective inference or post-selection inference; we will view rejection of $H_{0}^{1:p}$ as the “selection event”. The selective inference framework was laid out in Fithian et al. (2017) and further explored in a variety of settings (Taylor & Tibshirani, 2015; Lee et al., 2016; Zhang et al., 2019; Chen & Bien, 2020; Jewell et al., 2022; Chen & Gao, 2024; Chen et al., 2023; Chen & Witten, 2023; Gao et al., 2024; Neufeld et al., 2022, among others). Our goal also relates to the discussion of selective reporting of confidence intervals in Benjamini & Yekutieli (2005). Most related to our work is Heller et al. (2019), who considered inference following aggregate testing for a coefficient that follows a multivariate normal distribution with known covariance. We instead consider inference on the regression coefficients in Step 2 of Box 1.
@@
    We provide a valid end-to-end approach for conducting F-screening in the sense of Box 1. That is, we enable reporting of all of the standard multiple linear regression outputs—such as p-values, point estimates, and confidence intervals—conditional on rejection of $H_{0}^{1:p}$ in Step 1; we will refer to them as selective p-values, point estimates, and confidence intervals. Such a prospective analysis is suitable for a data analyst whose analysis plan is as in Box 1.
@@
    Furthermore, we show that the selective p-values can be computed using only the standard multiple linear regression outputs without access to the raw data. This is particularly critical because, while it is firmly established that selective reporting is common practice in the published literature (Chan et al., 2004), the raw data is typically not available to the statistician who wishes to correct for it.
@@
    We specialize the aforementioned approaches to the setting of analysis of variance (ANOVA), in which the F-screening procedure in Box 1, with Step 2 conducted using a “standard” approach that does not account for Step 1, is particularly pervasive (see, e.g., Fig. 1).
@@
Figure 1: A subset of Table 1 from Keil et al. (2023) in which F-screening was performed (with standard, and thus invalid, inference in Step 2 of Box 1). The $n=826$ participants in the dataset are split into three age groups: under 65 years old, between 65 and 85 years old, and over 85 years old. Each row in the table corresponds to a continuous variable of interest, with means and standard deviations reported for each age group. For each of these continuous variables, ANOVA is conducted to test for a difference in means between the three age groups. The associated p-value is reported in the “P Value” column. If this p-value is small, then post hoc pairwise tests are conducted, with p-values reported in the last column. The p-values in the last column are invalid since they do not account for the fact that they were computed only conditional on the fact that $H_{0}^{1:p}$ (2) was rejected.
@@
Addressing the inferential problems posed by F-screening is especially important from an applied perspective. A theoretical statistician unaccustomed to applied collaboration might believe that F-screening is an unusual procedure unlikely to be conducted in practice, and perhaps even that conducting Step 2 of Box 1 without accounting for Step 1 is so “obviously wrong” that no data analyst would consider it. The theoretical statistician would be mistaken on both counts. Statistical references that encourage F-screening with “standard” inference in Step 2 as part of a routine data analysis abound, and are authored by many of the luminaries in the field. [^1] Fisher (1935, §24) writes: “If the yields of the different varieties in the experiment fail to satisfy the test of significance they will not often need to be considered further . . . If, however . . . the null hypothesis has been falsified . . . we shall thereafter proceed to interpret the differences between the varietal yields . . . and shall be concerned to know with what precision these different yields have been evaluated.” A few years later, Lindquist (1940, page 98) writes “The student may well ask why it was necessary to apply the $F$-test at all, if we were subsequently going to test the individual differences anyway. The answer is that the $F$-test tells us whether it is worthwhile to test the individual difference at all. If the $F$-test had not proven significant, we would have known at once that all observed differences in . . . means could be due to chance alone. In that case it would not only have been unnecessary, but decidedly improper to apply the $t$-test to individual differences.” The abstract of Scheffé (1953) poses the following question: “Under the usual assumptions, if the conventional $F$-test of $H:\mu_{1}=\mu$ rejects $H$, what further inference are valid about the contrasts $H_{i}$?” Scheffé (1953) later remarks that “The problem of making further inferences about the contrasts, arising when the $F$-test rejects $H$, has been considered by various writers, including Fisher (1935), Newman (1939), Tukey (1949), Nandi (1951).” As pointed out in Olshen (1973), to the extent that these references consider selection, they focus on the multiplicity issues that arise from testing $H_{0}^{j}:\beta_{j}=0$, $j=1,\ldots,p$, rather than on the challenge that arises from inference conditional on rejecting $H_{0}^{1:p}$. See Appendix A for simulation results demonstrating that multiple testing corrections do not correct for selection in the F-screening procedure. Thus, it is hardly surprising that in the scientific literature, conducting Step 2 of Box 1 with a “standard” approach that does not account for having conducted Step 1 is often viewed as the “correct” analysis. [^2] As one of many possible examples, Webb et al. (2024) describe their analysis as follows: “A one-way ANOVA revealed that amygdala reactivity was not significantly different by trajectory groups . . . Therefore, additional analyses testing whether greenspace was associated with trajectory assignment via amygdala reactivity were not conducted.” In clinical journals, F-screening is a routine part of “Table 1,” in which patient characteristics are summarized; an example is shown in Fig. 1.
@@
Finally, we note that F-screening is almost certainly conducted much more frequently than the published literature suggests, both because (i) it is typically unknown whether an analysis would have been included in a published paper if $H_{0}^{1:p}$ had not been rejected, and (ii) an analysis for which $H_{0}^{1:p}$ is not rejected may not be published at all. Given that F-screening is both commonplace and problematic if inference in Step 2 does not account for Step 1, it came as a surprise to us that—to the best of our knowledge—no statistical solution has yet been proposed for valid inference in Step 2. The rest of this paper fills this gap.
@@
The paper is organized as follows. In Section 2, we propose a conditional selective inference approach for valid inference in Step 2 of Box 1, and we establish its theoretical properties in the case that the error variance is known; the “selection” event is rejection of $H_{0}^{1:p}$ in Step 1. Additionally, we propose a conditionally “debiased” estimator of the error variance that can be used in practice. In Section 3, we show that it is possible to “correct” the results of an analysis in which F-screening was conducted with “standard” (non-selective) inference in Step 2, without having access to the raw data. Section 4 specializes the results in Sections 2 and 3 to the case of ANOVA. Finally, Sections 5 and 6 contain simulation results and two re-analyses of published findings, respectively. We conclude with a discussion in Section 7. The main text focuses on developing valid p-values for Step 2 of Box 1; confidence intervals and point estimates are deferred to Appendix B. Proofs of all results are given in Appendix E.
@@ 2.1 Preliminaries
Let $P_{A}:=A(A^{\top}A)^{-1}A^{\top}$ denote the projection onto the column space of a matrix $A$ with full column rank, and let $A_{-S}$ denote the matrix $A$ with the columns indexed by $S\subseteq\{1,\ldots,p\}$ removed. Furthermore, let $\mathbbm{1}_{n}$ denote the $n$-dimensional column vector of 1’s, and $I_{n}$ the $n\times n$ identity matrix. This section restates some standard results in the theory of linear models that will prove useful in what follows (a more thorough treatment can be found elsewhere, e.g., Seber & Lee, 2003).
Consider $n$ independent observations from the model in (1). We assume that the intercept is not of scientific interest, and so both the initial test of $H_{0}^{1:p}$ and subsequent tests (if any) will operate on centered data. To do this, we let $U\in\mathbb{R}^{n\times(n-1)}$ be any matrix satisfying both $U^{\top}U=I_{n-1}$ and $UU^{\top}=I_{n}-P_{\mathbbm{1}_{n}}$, and then with slight abuse of notation we redefine $Y:=U^{\top}Y$, $X_{1}:=U^{\top}X_{1}$, and so on. Thus, in what follows, we will treat $\mathbb{R}^{n-1}$ as the ambient space and dispense with the intercept. For now, we assume that $\sigma^{2}$ is known (but that $F$-statistics are used nonetheless); we address the setting of unknown variance in Section 2.3.
First, consider the null hypothesis $H_{0}^{1:p}$ given in (2). The $F$-statistic defined as
@@
follows an $F_{p,n-p-1}$ distribution (that is, an $F$ distribution with $p$ and $n-p-1$ degrees of freedom) under $H_{0}^{1:p}$. Next, fix $H_{0}^{M}$ by letting $M\subseteq\{1,\ldots,p\}$ denote a (non-random) set of indices with cardinality $\lvert M\rvert=m$, and let $\beta_{M}$ denote the subset of coefficients indexed by $M$ (in the simplest case, $M=\{j\}$ for fixed $j\in\left\{1,\ldots,p\right\}$). Under the null hypothesis $H_{0}^{M}:\beta_{M}=0$, the $F$-statistic
@@
Fig. 2 displays the rejection regions corresponding to a test of $H_{0}^{1:p}$ using (3) and a test of $H_{0}^{M}$ using (4) in a simple setting with $n=3$ and $p=2$ and an orthogonal design matrix. The geometry of these two selection events interplay in non-trivial ways, and failing to account for the first when evaluating the second can yield invalid inference. By contrast, the selective test that we will develop in the next section will take the sequential nature of the analysis into account, and will adjust the geometry of the second test accordingly.

### v2

@@ Abstract
[…]
@@ 1 Introduction
at level $\alpha_{0}\in(0,1)$ in order to determine whether the model in (1) merits further study. If they fail to reject $H_{0}^{1:p}$, then (1) receives no further consideration, and the fact that (1) was considered may not even be reported. On the other hand, if (and only if) they reject $H_{0}^{1:p}$, they then proceed to test $H_{0}^{M}:\beta_{M}=0$ for fixed $M\subseteq\left\{1,\ldots,p\right\}$, where $\beta_{M}:=\{\beta_{j}:j\in M\}.$ Because $H_{0}^{1:p}$ is typically tested by means of an $F$-statistic, we refer to this procedure—summarized in Box 1—as F-screening. If conducted, the test in Step 2 of Box 1 is sometimes referred to as a post hoc test.
@@
Box 1: F-screening in the model (1). Step 1: Test $H_{0}^{1:p}$ as given in (2) at level $\alpha_{0}$, using an $F$-test. Step 2: Only if $H_{0}^{1:p}$ is rejected, conduct inference on $\beta_{M}$ for fixed $M\subseteq\{1,\ldots,p\}$.
@@
To see that conducting “standard” inference in Step 2 of F-screening will not provide the usual statistical guarantees, consider the extreme setting where $p=1$ and $\beta_{1}=0$. In this case, $H_{0}^{1:p}\equiv H_{0}^{1}$. Suppose we carry out F-screening in Box 1 using the “standard” $t$-test for $H_{0}^{1}$ in Step 2. Then for fixed $\alpha_{0}$, the tests in Steps 1 and 2 are identical, and so we will report a result in Step 2 only if $H_{0}^{1}$ is rejected. Thus the Type 1 error rate among reported results equals one! The left plot in Fig. 5 illustrates the inflation of Type 1 error that results from conducting “standard” inference in Step 2 in a setting with $p=10$ and $n=100$.
@@
As pointed out by Olshen (1973), conducting a test of $H_{0}^{M}$ that accounts for rejection of $H_{0}^{1:p}$ is distinct from the task of correcting multiple tests of the form $H_{0}^{M}:\beta_{M}=0$ for multiple comparisons. The latter has been very well-explored in the literature: see Dunn (1961); Šidák (1967); Tukey (1949); Scheffé (1953). For instance, if our interest in Step 2 of Box 1 lies in $H_{0}^{j}:\beta_{j}=0$ for $j\in\{1,\ldots,p\}$, then we could consider testing each null hypothesis using a standard test, and applying a Bonferroni correction to account for the fact that $p$ tests were conducted. However, this will not address the fact that F-screening was performed: namely, that we test $\{H_{0}^{j}\}_{j=1}^{p}$ if and only if we have rejected $H_{0}^{1:p}$. In Appendix A, we elaborate on how the guarantees provided by multiple testing corrections—such as family-wise error rate control—differ from those achieved by our proposed method. Further, we describe a closed-testing procedure that parallels F-screening and show that its guarantees are quite different. Fig. A.1 demonstrates empirically that existing approaches fail to provide valid inference conditional on the rejection of $H_{0}^{1:p}$.
@@
Our goal of conducting inference on $\beta_{M}$ that is valid conditional on the fact that we rejected $H_{0}^{1:p}$ falls under the umbrella of selective inference or post-selection inference; we will view rejection of $H_{0}^{1:p}$ as the “selection event.” The selective inference framework was laid out in Fithian et al. (2017) and further explored in a variety of settings (Taylor & Tibshirani, 2015; Lee et al., 2016; Zhang et al., 2019; Chen & Bien, 2020; Jewell et al., 2022; Chen & Gao, 2024; Chen et al., 2023; Chen & Witten, 2023; Gao et al., 2024; Neufeld et al., 2022, among others). Our goal also relates to the discussion of selective reporting of confidence intervals in Benjamini & Yekutieli (2005). Most related to our work is Heller et al. (2019), who considered inference following aggregate testing for a coefficient that follows a multivariate normal distribution with known covariance. We instead consider inference on the regression coefficients in Step 2 of Box 1.
@@
    We provide a valid end-to-end approach for conducting F-screening in the sense of Box 1. That is, we enable reporting of all of the standard multiple linear regression outputs—such as p-values, point estimates, and confidence intervals—conditional on rejection of $H_{0}^{1:p}$ in Step 1; we will refer to them as selective p-values, point estimates, and confidence intervals.
@@
    We establish consistency and selective Type 1 error control (the Type 1 error conditional on rejection of $H_{0}^{1:p}$) of the test resulting from our selective p-value.
@@
    We characterize the Fisher information about $\beta_{M}$ resulting from our procedure, and show that it compares favorably to that of a procedure that relies on sample splitting.
  4. 4.
    We show that the selective p-values can be computed using only the standard multiple linear regression outputs without access to the raw data. This is particularly critical because, while it is firmly established that selective reporting is common practice in the published literature (Chan et al., 2004), the raw data is typically not available to the statistician who wishes to correct for it.
  5. 5.
@@
    We specialize our approach to the setting of analysis of variance (ANOVA), in which the F-screening procedure in Box 1, with Step 2 conducted using a “standard” approach that does not account for Step 1, is particularly pervasive (see, e.g., Fig. 1).
@@
Figure 1: A subset of Table 1 from Keil et al. (2023). The $n=826$ participants in the dataset are split into three age groups: under 65 years old, between 65 and 85 years old, and over 85 years old. Each row in the table corresponds to a continuous variable of interest, with means and standard deviations reported for each age group. For each of these continuous variables, ANOVA is conducted to test for a difference in means between the three age groups (corresponding to Step 1 of Box 1). The associated p-value is reported in the “P Value” column. The last column holds p-values for “post hoc tests” which are typically conducted only if the p-value associated with the ANOVA test is small (corresponding to Step 2 of Box 1). In this case, the p-values in the last column would be invalid since they do not account for the fact that they were only computed because $H_{0}^{1:p}$ (2) was rejected.
@@
Addressing the inferential problems posed by F-screening is especially important from an applied perspective. A theoretical statistician unaccustomed to applied collaboration might believe that F-screening is an unusual procedure unlikely to be conducted in practice, or that conducting Step 2 of Box 1 without accounting for Step 1 is “obviously wrong” and no data analyst would consider it. The theoretical statistician would be mistaken on both counts. Statistical references that encourage F-screening with “standard” inference in Step 2 as part of a routine data analysis abound, and are authored by many of the luminaries in the field. [^1] Fisher (1935, §24) writes: “If the yields of the different varieties in the experiment fail to satisfy the test of significance they will not often need to be considered further . . . If, however . . . the null hypothesis has been falsified . . . we shall thereafter proceed to interpret the differences between the varietal yields . . . and shall be concerned to know with what precision these different yields have been evaluated.” A few years later, Lindquist (1940, page 98) writes “The student may well ask why it was necessary to apply the $F$-test at all, if we were subsequently going to test the individual differences anyway. The answer is that the $F$-test tells us whether it is worthwhile to test the individual difference at all. If the $F$-test had not proven significant, we would have known at once that all observed differences in . . . means could be due to chance alone. In that case it would not only have been unnecessary, but decidedly improper to apply the $t$-test to individual differences.” The abstract of Scheffé (1953) poses the following question: “Under the usual assumptions, if the conventional $F$-test of $H:\mu_{1}=\mu$ . . . rejects $H$, what further inference are valid about the contrasts $H_{i}$?” Scheffé (1953) later remarks that “The problem of making further inferences about the contrasts, arising when the $F$-test rejects $H$, has been considered by various writers, including Fisher (1935), Newman (1939), Tukey (1949), Nandi (1951).” As pointed out in Olshen (1973), to the extent that these references consider selection, they focus on the multiplicity issues that arise from testing $H_{0}^{j}:\beta_{j}=0$, $j=1,\ldots,p$, rather than on the challenge that arises from inference conditional on rejecting $H_{0}^{1:p}$. See Appendix A for simulation results demonstrating that multiple testing corrections do not correct for selection in the F-screening procedure. Thus, it is hardly surprising that in the scientific literature, conducting Step 2 of Box 1 with a “standard” approach that does not account for having conducted Step 1 is often viewed as the correct analysis. [^2] As one of many possible examples, Webb et al. (2024) describe their analysis as follows: “A one-way ANOVA revealed that amygdala reactivity was not significantly different by trajectory groups . . . Therefore, additional analyses testing whether greenspace was associated with trajectory assignment via amygdala reactivity were not conducted.” In clinical journals, F-screening is a routine part of “Table 1,” in which patient characteristics are summarized; an example is shown in Fig. 1.
@@
Finally, we note that F-screening is likely conducted much more frequently than the published literature suggests, both because (i) it is typically unknown whether an analysis would have been included in a published paper if $H_{0}^{1:p}$ had not been rejected, and (ii) an analysis for which $H_{0}^{1:p}$ is not rejected may not be published at all. Given that F-screening is both commonplace and problematic if inference in Step 2 does not account for Step 1, it came as a surprise to us that—to the best of our knowledge—no statistical solution has yet been proposed for valid inference in Step 2. The rest of this paper fills this gap.
@@
The paper is organized as follows. In Section 2, we propose a conditional selective inference approach for valid inference in Step 2 of Box 1. We establish its theoretical properties and derive an expression for the leftover Fisher information after rejection of $H_{0}^{1:p}$ (2) in Step 1. We also propose an estimator of $\sigma^{2}$ that is unbiased and consistent conditional on rejection of $H_{0}^{1:p}$, and that yields valid p-values. Section 3 shows that it is possible to compute selective p-values using the results of a “standard” (non-selective) analysis in Step 2, without having access to the raw data. Section 4 specializes the results in Sections 2 and 3 to the case of ANOVA. Simulation results are in Section 5 (and Appendix C), and re-analyses of published findings in Section 6. The discussion is in Section 7. The main text focuses on developing valid selective p-values; confidence intervals and point estimates are deferred to Appendix B. Appendix D contains mathematical details for Figures 2 and 3. Proofs of all results are given in Appendix E.
@@ 2.1 Preliminaries
Let $P_{A}:=A(A^{\top}A)^{-1}A^{\top}$ denote the projection onto the column space of a matrix $A$ with full column rank, and let $A_{-S}$ denote the matrix $A$ with the columns indexed by $S\subseteq\{1,\ldots,p\}$ removed. Furthermore, let $\mathbbm{1}_{n}$ denote the $n$-dimensional column vector of 1’s, and $I_{n}$ the $n\times n$ identity matrix. This section restates some standard results in the theory of linear models (a more thorough treatment can be found elsewhere, e.g., Seber & Lee, 2003).
Consider $n$ independent observations from the model in (1). We assume that the intercept is not of scientific interest, and so both $H_{0}^{1:p}$ and $H_{0}^{M}$ will operate on centered data. To do this, we let $U\in\mathbb{R}^{n\times(n-1)}$ be any matrix satisfying both $U^{\top}U=I_{n-1}$ and $UU^{\top}=I_{n}-P_{\mathbbm{1}_{n}}$, and then with slight abuse of notation we redefine $Y:=U^{\top}Y$, $X_{1}:=U^{\top}X_{1}$, and so on. Thus, in what follows, we will treat $\mathbb{R}^{n-1}$ as the ambient space and dispense with the intercept. For now, we assume that $\sigma^{2}$ is known (but that $F$-statistics are used nonetheless); we address the setting of unknown variance in Section 2.4.
First, consider the null hypothesis $H_{0}^{1:p}$ given in (2). Under $H_{0}^{1:p}$, the $F$-statistic defined as
@@
follows an $F_{p,n-p-1}$ distribution (that is, an $F$ distribution with $p$ and np1n-p-1 degrees of freedom). Next, fix $H_{0}^{M}$ by letting $M\subseteq\{1,\ldots,p\}$ denote a set of indices with cardinality $\lvert M\rvert=m$, and let $\beta_{M}$ denote the subset of coefficients indexed by $M$ (in the simplest case, $M=\{j\}$ for fixed $j\in\left\{1,\ldots,p\right\}$). Under the null hypothesis $H_{0}^{M}$, the $F$-statistic
@@
Fig. 2 displays the rejection regions corresponding to a test of $H_{0}^{1:p}$ using (3) and a test of $H_{0}^{M}$ using (4) in a simple setting with $n=3$, $p=2$, and $X$ orthogonal. The geometry of these two events interplay in non-trivial ways, and failing to account for the first when evaluating the second can yield invalid inference. By contrast, the selective test that we will develop in the next section will take the sequential nature of the analysis into account, and it will adjust the geometry of the second test accordingly.